### v3

# Symmetry Breaking in Chemical Systems: Engineering Complexity through Self-Organization and Marangoni Flows

Abstract
Far from equilibrium, chemical and biological systems can form complex patterns and waves through reaction-diffusion coupling. Fluid motion often tends to disrupt these self-organized concentration patterns. In this study, we investigate the influence of Marangoni-driven flows inside a thin layer of fluid ascending the outer surfaces of hydrophilic obstacles on the spatio-temporal dynamics of chemical waves in the modified Belousov-Zhabotinsky reaction. Our observations reveal that circular waves originate nearly simultaneously at the obstacles and propagate outward. In a covered setup, where evaporation is minimal, the wavefronts maintain their circular shape. However, in an uncovered setup with significant evaporative cooling, the interplay between surface tension-driven Marangoni flows and gravity destabilizes the wavefronts, creating distinctive flower-like patterns around the obstacles. Our experiments further show that the number of petals formed increases linearly with the obstacle’s diameter, though a minimum diameter is required for these instabilities to appear. These findings demonstrate the potential to ’engineer’ specific wave patterns, offering a method to control and direct reaction dynamics. This capability is especially important for developing microfluidic devices requiring precise control over chemical wave propagation.
keywords:
Belousov-Zhabotinsky reaction, Marangoni flows, Chemo-hydrodynamic patterns, Hydrodynamic fingering instabilty, Reaction-diffusion systems
Sangram Gore, Binaya Paudyal, Mohamed Ali, Nader Masmoudi, Albert Bae, Oliver Steinbock, Azam Gholami*
[…]
{affiliations}
Dr. S. Gore, B. Paudyal, Dr. M. Ali, Prof. N. Masmoudi, Prof. A. Gholami
Science Division, New York University Abu Dhabi, Abu Dhabi, UAE
Email Address: azam.gholami@nyu.edu
Prof. N. Masmoudi
Courant Institute of Mathematical Sciences, New York University, New York, USA
Prof. A. Bae
Lewis &\&& Clark College, Portland, Oregon, USA
Prof. O. Steinbock
Department of Chemistry and Biochemistry, Florida State University, Tallahassee, Florida, USA
1 Introduction
Symmetry breaking is a fundamental concept in physics, biology and chemistry where small perturbations can lead to the emergence of complex structures and patterns from initially uniform states [1, 2]. In fluid dynamics, this phenomenon is exemplified by the transition from simple, laminar flows to intricate turbulent structures, driven by instabilities that disrupt the symmetry of the system. In biology, Alan Turing’s reaction-diffusion theory proposes that the initial symmetry in embryos can be broken by the interaction of two diffusible molecules, whose dynamic interplay leads to the formation of complex patterns [3]. Similarly, in chemical systems like the Belousov-Zhabotinsky (BZ) reaction, symmetry breaking can give rise to a rich spectrum of spatio-temporal structures, such as rotating spirals and target waves as well as Turing patterns [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23]. These patterns, resulting from the interplay between nonlinear reaction kinetics and diffusion, offer profound insights into the mechanisms of self-organization in both natural and synthetic systems [1, 24, 25].
The interaction between chemical reactions and convective fluid motion, particularly through mechanisms such as Marangoni-driven flows, profoundly influences the formation and breakdown of complex spatio-temporal patterns [26, 27, 28, 29]. In pioneering studies on the BZ reaction [30, 31, 32, 33, 34], it was observed that in confined environments, chemical instabilities manifest themselves as propagating waves. In contrast, in systems with a free surface, the presence of Marangoni flows lead instead to the formation of cellular structures [35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 12]. These flows are initiated by gradients in surface tension [46, 44], which might arise from either temperature variations or differences in solute concentrations across the fluid’s surface, causing the fluid to flow from regions of lower to higher surface tension. For instance, in the ferroin-catalyzed BZ reaction, Fe2+ complex has a higher surface activity than Fe3+ complex which is attributed to the differing hydrophilic properties of the catalyst’s oxidized (more hydrophilic) and reduced forms. This difference leads to an increase in surface tension of approximately 2 mN/m when the system transitions from the reduced to the oxidized state [46]. In an uncovered setup, where evaporative cooling is significant, the interplay of chemical waves, buoyancy-driven flows due to density gradients, and Marangoni flows due to surface tension gradients collectively influences the resulting chemo-hydrodynamic patterns [31, 47, 48, 41, 40, 49, 34, 50].
Evaporation-induced Marangoni flows can be intensified by introducing hydrophilic vertical or inclined obstacles within the reactive medium. The thin film rising along the outer surfaces of these obstacles evaporates more quickly than the bulk fluid, creating thermal/solutal surface tension gradients. The delicate balance between upward-driving surface tension forces and gravity-induced downward motion leads to undulations at the fluid interface, potentially developing into finger-like patterns if the fluid becomes sufficiently unstable [51, 52, 53, 54, 55, 56, 57]. The obstacle’s geometry also affects the dynamics at the contact line region of thin films. A cylindrical obstacle introduces unique conditions, such as azimuthal curvature and periodic boundary conditions, absent in flat obstacles [57]. This periodicity ensures that an integer number of fingers can form along a cylinder, a condition that is not required for a flat obstacle, and the smallest number of fingers that can form along the contact line of a cylinder is one [57].
Building on these groundworks, our study presents a novel approach to explore the intricate interaction between evaporation driven Marangoni flows and chemical waves in the modified BZ reaction in the presence of hydrophilic obstacles. We observe that in a covered experimental setup with minimal evaporative cooling, the obstacles act as focal points that consistently emit synchronous circular waves and maintain their shape over multiple cycles. In contrast, thermal/solutal driven Marangoni flows in an uncovered setup, where evaporative cooling is more pronounced, trigger a fingering instability that leads to the formation of distinct flower-like wave patterns. Significantly, the number of petals increases with the diameter of the obstacles, and a minimum threshold value for the diameter is required for this instability to occur. We performed numerical simulations of a reaction-diffusion model to explain the observed wave dynamics and gain deeper insights into the underlying mechanisms.
2 Results
2.1 Experimental Results
Our quasi 2D setup consists of a periodic array of millimeter-sized PDMS-based cylindrical obstacles (see Figure 1A) that influence the spatio-temporal dynamics of chemical waves in the modified BZ reaction (see Method Section). This version of the BZ reaction, known as CHD-BZ, employs 1,4-cyclohexanedione as a bubble-free organic substrate and ferroin as a high-absorbance redox indicator [58, 59, 60, 61, 62, 63, 64, 65]. The addition of the ferroin redox indicator to the reaction mixture renders the oscillations visible, with periodic color changes between blue and red. The PDMS has undergone plasma treatment to make it hydrophilic, which allows the fluid to climb up around these obstacles, as illustrated in Figure 1B. We observe circular oxidation waves in blue color (high value of ferriin concentration Fe3+) that initiate almost synchronously at the pillars and propagate inside the reduced state in red color (high concentration of ferroin Fe2+) at a speed of approximately 7 mm/min and recurring every 45 sec (Figure 1C-L). When these waves collide, they mutually annihilate each other. The time evolution of these waves is strongly influenced by whether the setup is covered or uncovered (as illustrated in Figure 2B), which impacts the evaporation rate of the liquid (Figure 2C). Below, we discuss the results for both setups.
Figure 1: Wave dynamics in a covered versus an uncovered setup. (A) A standard experimental setup featuring a periodic arrangement of PDMS-based cylindrical obstacles designed to fit within a 10 cm Petri dish. (B) A side view depicting the fluid profile around an exemplary hydrophilic pillar. (C-D) Top views of concentric chemical waves in a covered setup, captured at two distinct times, show the waves initiating almost simultaneously at the pillars and spreading outward. The waves retain their circular shapes over time. (E-F) Top views of chemical waves captured at two distinct times in an uncovered setup: initially, concentric waves form around the obstacles (E), and later, the wavefronts fracture to create flower-like patterns (F). (G) A space-time plot along the orange line indicated in (C) demonstrates that the pillars serve as the wave centers. Black bars show the position of the obstacles. (H) A space-time plot traced along the cyan-colored circle shown in (C) illustrates that the circular waves retain their shape through multiple wave cycles. (I) The space-time plot along the yellow line shown in (E). (J) The space-time diagram along the green circle indicated in (E) illustrates that after two cycles of circular waves, the wavefronts break. In this experiment, where pillars are 1 mm in diameter, 3 mm tall, and spaced 5 mm apart, most pillars develop three petals. (K-L) Images displaying the entire Petri dish from two separate experiments: one covered and the other uncovered. The color-coded wave period around each pillar is also shown in seconds. The timestamps reflect the elapsed time since the start of recording, which began shortly after the chemical solution was introduced into the setup. These images are typical of at least 50 experiments.
2.1.1 Covered setup: Circular waves initiate synchronously at the pillars and maintain their circular shape
In a covered setup, chemical waves begin nearly simultaneously on the pillars, with a wave period of about 45 sec, and maintain their circular shape across multiple wave cycles (Figure 1C-D and Video 1). A space-time plot along an arbitrary line crossing the obstacles illustrates that they serve as wave centers (Figure 1G). Additionally, to visualize the wave dynamics over several cycles, we compiled the light intensity data along a circle surrounding an arbitrary pillar, confirming that the wavefronts consistently retain their circular form (Figure 1H). Figure 1K and Video 2 present a separate experiment where the entire Petri dish is imaged, demonstrating similar wave dynamics. Fourier analysis of the waves emanating from the obstacles indicates that the majority have a periodicity of 45 sec (see Figure S1A-D).
2.1.2 Uncovered setup: Circular waves begin synchronously at the pillars and eventually evolve into flower-like patterns
Figure 2: Number of petals depends on the pillar diameter and can vary over time. (A) A PDMS mold with alternating pillar diameters of 1 mm and 1.5 mm is shown. (B) A sketch illustrates the covered versus uncovered experimental setup. (C) In the uncovered setup with evaporative cooling, the evaporation rate is approximately 2.5$\times$10-3 mg/(min mm2), which is about seven times higher than in the covered setup. (D-E) Top views of chemical waves in an uncovered setup captured before and after the onset of instability in an experiment where pillar diameters alternate between rows, with 1.5 mm pillars in the top row and 1 mm pillars in the second row. Pillars with a diameter of 1.5 mm generally develop four petals, whereas those with a diameter of 1 mm typically produce three petals. (F-H) Space-time plots along the highlighted circles in panel (D) illustrate the progression of the instability. (I) An example of a pillar that initially developed three petals, but over time, two of the petals merged.
In an uncovered setup, where evaporation is more significant (Figure 2C), the obstacles continue to act as centers that emit synchronous concentric waves (Figure 1E, I and Video 3). However, after approximately two cycles of these circular waves, an instability in the wavefront becomes apparent: the circular wavefronts fragment, leading to the formation of striking flower-like patterns (see Figure 1F, J). Figure 1L and Video 4 present a separate experiment where the entire Petri dish is imaged, demonstrating similar wavefront instability. Consistent with the covered setup, Fourier analysis shows that the waves originating from the obstacles typically have a periodicity of about 45 seconds (see Figure S1E-H).
Our experimental findings indicate that the number of petals is influenced by the diameter of the pillars. Typically, pillars with a diameter of 1 mm form three petals (Figure 1F), while those with a diameter of 1.5 mm often produce four petals (Figure 2 and 3), and pillars measuring 3.8 mm in diameter generate six to seven petals (Figure 4 and Video 5). Figure 2 presents an illustrative experiment where the pillar diameters alternate between 1 mm and 1.5 mm across neighboring rows, with the top row featuring pillars of 1.5 mm in diameter. In this setup, the 1 mm pillars typically form patterns with three petals, whereas the 1.5 mm pillars display flower patterns consisting of four petals (Figure 2E-I and Video 6).
Interestingly, for a small subset of the pillars, we observed that the number of petals can vary over time, as illustrated in Figure 2I. In a related observation, Figure 3 depicts an experiment where, after two cycles of circular wavefronts, the wavefronts around the pillars break. Initially, four petals form, but as time progresses, a fifth petal develops (Figure 3C). The synchronization between the waves initiated at different pillars is demonstrated through a phase map calculated using the Hilbert transform (see Method Section), as shown in Figure 3A-B. In this experiment, where we have captured images of the entire small Petri dish, the waves initiated at the PDMS boundary also exhibit wavefront instability, similar to those observed at the pillars (see Figure 3D and Video 7).
Figure 3: Wavefront instability at the boundary in an uncovered setup. (A-B) Pillars with a diameter of 1.85 mm typically form four petals. The color-coded phase map, generated using Hilbert transform, indicates that the pillars simultaneously act as wave centers. (C) A space-time plot along the cyan-highlighted circle in panel (A) shows the development of the instability, with a noticeable evolution in the number of petals over time. The wave period is 50 sec. (D) A space-time plot along the orange-highlighted circle demonstrates the wavefront instability that also manifests at the PDMS wall on the periphery. This experiment is conducted in a small Petri dish.
Figure 4: Effect of the pillar diameter on the number of petals. (A-D) As the pillar diameter decreases, so does the number of petals formed. For instance, pillars with a diameter of 3.8 mm typically develop around 7 petals, while those with a diameter of 2.8 mm produce about 6 petals. This number drops to 4 petals for pillars with a diameter of 1.8 mm, and further reduces to 3 petals for those with a diameter of 1 mm. (E) Pillars with a diameter of 0.5 mm exhibit no wavefront instability, which verifies that there is a minimum pillar perimeter required for such instability to develop. (F) The number of petals increases linearly with the diameter of the pillar. Below a critical diameter, wavefronts do not break and remain circular.
To unravel the mechanisms (i) governing the formation of synchronous circular waves around the pillars and (ii) contributing to the instability of the wavefront in an uncovered setup, we began by examining the hypothesis that the permeability of PDMS to bromine gas (and potentially other chemicals) facilitates wave formation on the pillars, an idea supported by literature [66, 67, 68]. To test this possibility, we conducted experiments using PDMS pillars coated with a thin layer of gold (approximately 200 nm thick, see Figure S2A). Contrary to expectations, the pillars continued to function as centers for synchronous wave activity even with the gold coating, which blocks chemical absorption (Figure S2B and Video 8). Additionally, we conducted experiments using obstacles made of acrylic (PMMA) with plasma treatment and observed that these pillars also functioned as centers for wave activity (Figure 5A-C). These results confirm that the absorption of chemicals into PDMS is not the primary mechanism driving the formation of concentric waves around the obstacles.
Next, we explored the impact of the thin liquid layer ascending the PDMS-made pillars, which are hydrophilic due to plasma treatment. Our investigation involved several experimental approaches. Initially, we conducted experiments with pillars that had not undergone plasma treatment and observed that most of the pillars did not act as centers for wave generation (see Figure 6A-B and Video 9). Subsequently, we decreased the height of the pillars from 3 mm to 1 mm, ensuring that these shorter pillars were submerged to the same fluid level. We noted that the submerged pillars also did not act as wave centers (Video 10).
Finally, to assess whether the formation of circular waves around the pillars was influenced by the proximity of adjacent pillars, we conducted experiments with a single isolated pillar. As shown in Figure 7 and Video 11, this solitary pillar also acted as a wave center and exhibited wavefront instability in an uncovered setup. Notably, under uncovered conditions, only waves originating from the pillar and the PDMS wall at the periphery exhibited wavefront instability, while off-center waves retained their circular shape without fragmentation. These findings collectively highlight the crucial role of the climbing liquid layer on the pillars in both the generation of synchronous circular waves and the onset of wavefront instability.
Figure 5: Experiments with PMMA-made obstacles. (A) Side view of the cylindrical obstacle array made from PMMA, with walls forming a square boundary. The fluid climbs up the plasma-treated hydrophilic obstacles as well as the wall boundary. (B) Images are subtracted every 3 seconds to better visualize the wavefronts initiating in the thin film region, wetting the pillars and the walls, and traveling downward. The obstacles act as wave centers, as shown in the space-time plot in the red box, which stacks up light intensity along the red line. (C) Initially, the wavefronts are planar along the wall and circular around the pillars, but they rapidly deform, developing into finger-like patterns. This wavefront instability on the wall is clearly visible in the space-time plot in the green box, created by stacking up the light intensity along the green line on the wall boundary. (D) A top view of the waves around the obstacles, illustrating the breaking of the wavefront and the emergence of flower-like patterns. These images are representative of at least 10 experiments.
Figure 6: Hydrophilic versus hydrophobic obstacles: An overview of a partitioned, uncovered Petri dish taken at two distinct times: before (A) and after (B) the emergence of instability. The upper portion of the dish has been plasma-treated to render PDMS hydrophilic, whereas the lower section has been left untreated, maintaining its hydrophobic properties (see Video 9). A zoomed-in side view of the fluid near the pillars is also presented, emphasizing the rising fluid around the hydrophilic pillar. (C-E) Side-view imaging of the fluid climbing the hydrophilic pillars confirmed that the contact line remains pinned. As the liquid evaporates, the contact angle gradually decreases (see Video 12). The cyan curves represent the air-liquid interface at $t$$=$0.
2.1.3 Experimental evidence of a fingering instability
Figure 7: A single pillar also acts as the wave center. (A-C) An isolated pillar serves as the wave center and exhibits wavefront instability in an uncovered Petri dish. Additionally, we observe wavefront instability at the PDMS walls around the periphery. (D) shows a flower-like pattern with six or seven petals at different time points that formed around the solitary central pillar.
Figure 8: Experiments with rounded rectangular obstacles and triangular obstacles. (A) In an experiment examining the impact of evaporation, PDMS-made obstacles with a rectangular body and rounded ends were used. (B) The sequence of wave pattern of the highlighted pillar in panel (A) indicates that instability mainly arises at the curved ends. (C) Oscillations of light intensity averaged over the red box shown in panel (B). (D) A space-time plot along the circular path marked in cyan in panel (B) illustrates the progression of wavefront instability. (E) A side view of an experiment featuring triangular PMMA obstacles within a square boundary, where waves predominantly initiate at the sharp edges. Space-time plots along the green and orange lines are also shown. The intensity averaged over the red box shows oscillations of 60 seconds. Notably, the chemical fingers are clearly visible along the wall.
As previously discussed, our experiments reveal that the climbing film on the external surfaces of cylindrical pillars is essential for the formation of synchronous circular waves and the onset of wavefront instability. We observe that wavefront instability predominantly occurs in uncovered experimental setups, where evaporation is more significant compared to covered setups. Notably, in our control experiments, where the CHD-BZ solution is covered with a thin layer of hexane to suppress evaporation, the obstacles continue to function as wave centers, but no wavefront instability is observed (Figure S2D-E).
Moreover, side-view imaging of the fluid climbing around the pillars confirms that the contact line remains pinned (Figure 6C-E and Video 12), indicating the presence of upward capillary flows that counteract fluid loss caused by evaporation. Within this ascending liquid layer, the thinner downstream precursor film is expected to evaporate faster than the thicker upstream film, creating slight temperature and/or concentration gradients. These gradients generate surface tension differences, driving Marangoni flows that ultimately disrupt the wavefront. Our hypothesis that waves originate from the thin ascending fluid layer is supported by our observations in Figure 5, using a PMMA mold with square wall boundaries. The side view of the waves confirms that they initiate at the thin film in the contact region, which wets the square wall boundaries, and then travel downward (Video 13). Importantly, in an uncovered setup, the downward-traveling wave along the wall boundary (as well as the waves initiated around the cylindrical obstacles) initially maintain a planar (circular) form but eventually break into fingers at later times (Figure 5B-C and D). This fingering instability is similarly visible in Figure 8E and Video 14 with PMMA-based triangular obstacles within square wall boundaries. Lastly, we note that the initiation of waves at the ascending thin film is consistent with a separate experiment in which the Petri dish was slightly tilted [37]. In this setup, we observed that synchronization waves initially appeared in regions with lower fluid thickness and subsequently propagated toward areas with higher fluid thickness (see Figure S5A-B and Video 15).
Literature reveals that Marangoni-driven thin films on planar or curved substrates exhibit a complex array of dynamics [57, 54, 69], dependent on the flat upstream film thickness trailing the advancing front. When the upstream film thickness is set by the competition effect of meniscus curvature and the surface tension gradient, the film is sufficiently thin for the advancing front to form a stationary compressive or Lax shock, which are well known to be unstable to contact line perturbations [69, 54, 57, 55]. It is important to note that surface tension gradients, which drive Marangoni flows, can result from temperature gradients, concentration gradients, or a combination of both. Concentration gradients often develop due to the evaporation of more volatile chemical components within the ascending thin film, similar to the mechanism observed in the tears of wine phenomenon [70]. The interplay between Marangoni flows and gravity can disrupt the contact lines, evolving into a finger pattern. Specifically, the linear stability analysis in Ref. [57] correlates linearly the number of fingers with the cylindrical pillar’s perimeter, aligning with our findings (see Figure 4). This analysis also suggests a minimum pillar diameter threshold below which fingering instability does not occur. To validate this, we reduced the pillar diameter to 0.5 mm in our uncovered setup experiments, observing that the wavefronts remained nearly circular (see Figure 4E and Figure S3).
Moreover, we tested the impact of evaporation on fingering instability using obstacles with a rectangular geometry and circular ends (see Figure 8A-D and Video 16). Given the increased surface area at the curved boundaries, we anticipated a higher evaporation rate here, which likely triggers the fingering instability predominantly at these locations, a hypothesis supported by our observations. A similar pattern was noted in our experiments with obstacles of triangular cross-section, where waves primarily started at the sharp boundaries (Figure 8E and Video 14). Additionally, we conducted experiments with both inverted tapered and tapered pillars, confirming that as long as a thin wetting layer surrounds the obstacles, they act as synchronous wave centers (Figure 9 and Videos 17-20). In the experiments with tapered pillars, we observed that wavefront instability first appeared in non-tapered pillars (our standard straight pillars with a 0 taper angle) and, after a time delay, in the tapered pillars as the angle gradually increased from 0 to approximately 26 (Figure 10). We attribute this effect to the higher evaporation rate in standard pillars compared to tapered pillars, due to their greater available surface area at the fluid-pillar contact line. Interestingly, we found that the average number of petals remains unchanged across different taper angles. However, as the perimeter of the pillars at the fluid-pillar contact line decreases with larger taper angles (see yellow circles in Figure 10B), the wavelength also decreases as the taper angle increases, ensuring that the same number of petals is accommodated.
Figure 9: Experiments with both inverted tapered and tapered pillars. (A) Conical pillars with base diameters ranging from 0.5 mm to 2 mm, with the top row displaying standard cylindrical pillars of 3 mm diameter. (B) Side view of centered waves observed before and after the onset of instability. The space-time within the green box represents the light intensity along the green line over time, revealing that the waves originate from the climbing thin film and propagate downward. (C-D) A close-up side view of two neighboring cones with a base diameter of 1 mm. The cyan curve illustrates the air-liquid interface at $t$$=$0. (E) Top view of the centered waves prior to the instability. (F) Tapered pillars with a 3 mm base diameter that gradually decrease to smaller diameters toward the top. (G-J) The side and top views of the centered waves are shown, along with the contact angle of the CHD-BZ solution around the hydrophilic conical structures at $t$$=$0. Over time, as evaporation progresses, the contact angle decreases ($t$$=$300 sec). The cyan curves in panel (H) represent the liquid-air interface at $t$$=$0; see Videos 17-20.
Figure 10: (A) Side-view of tapered pillars used for data processing of waves centered around them in panels (B–D), with the last pillar at the bottom representing the standard non-tapered pillar. (B-D) The top view highlights the circles along which the space-time plot in panel (C) is generated. The colored squares indicate the areas where light intensity is averaged and plotted over time in panel (D). Both the space-time plot and light intensity oscillations show that after the first cycle of nearly synchronous circular waves, the second round of waves with a broken wavefront first appears in the non-tapered pillar at the bottom, followed by a time delay in the tapered pillars. The two tapered pillars at the top, both tapered at approximately 26, exhibit the longest delay. The orange lines serve as visual guides (see Video 18).
Furthermore, we observed that during multiple cycles of covering and uncovering the setup, the wavefronts, which exhibit flower-like patterns when uncovered, revert to their circular form once the setup is covered, and vice versa (Videos 21-22). To further investigate the role of boundaries, we conducted a control experiment using an uncovered, plasma-treated glass Petri dish without any obstacles (Figure S4 and Video 23). Interestingly, we observed wavefront instability at the glass boundary, closely resembling the phenomena observed at the PDMS wall in Figure 3 and the PMMA wall in Figures 5 and 8E. A similar control experiment was performed in an uncovered Petri dish with a PDMS substrate (no pillars) and a PDMS wall boundary, where wavefront instability was only observed at the PDMS wall (see Video 24). Overall, these experiments confirm that the thin layer of climbing fluid at the boundary is critical for initiating waves, and that evaporation significantly contributes to the wavefront instability.
Our final experiment that clearly proves the important role of evaporation in spontaneous breaking of azimuthal-symmetry is an evaporating CHD-BZ droplet. The preferential evaporation of more volatile chemical components at the rim can contribute to the segregation of other chemicals near the rim and eventually break the azimuthal symmetry. Figure 11A-D and Video 25 show spot formation near the boundary of a droplet. A similar effect has been reported in water/1,2-hexanediol binary drops in Refs. [71, 72, 73, 74], where 1,2-hexanediol micro-droplets nucleate at the rim of the droplet. We note that in our evaporating droplet experiments, where micron-sized beads were added as tracer particles to track the flow, we observed the coffee-stain effect (see Video 26 and Figure 11E-H.). This effect is characterized by the formation of a ring-like structure of deposited beads after the droplet evaporates. The phenomenon arises from a combination of the non-uniform evaporation rate along the droplet interface and the presence of a pinned contact line, which generates a continuous capillary flow from the center of the droplet toward the contact line to compensate for liquid loss during evaporation [75, 76]. Furthermore, we identified a secondary instability at the droplet boundary (Figure 11H), which requires further investigation in future studies.
Figure 11: Symmetry breaking in an evaporating CHD-BZ droplet and coffee stain effect. (A-D) A droplet of volume $2.5~{}\mu l$ is deposited on a hydrophobic glass substrate. After about 500 sec, the azimuthal symmetry breaks and spots form at the rim evolving over time ($T$$=$22C, RH$=$$40\%$, $\theta_{0}$$=$12). The images are captured in the reflection mode of a confocal microscope; see Video 25. (E-H) In a separate experiment imaging the entire droplet, micron-sized beads were introduced to track the flow. During the first 5 minutes of evaporation, the droplet boundary remains pinned to the substrate, leading to the accumulation of tracer particles at the rim due to capillary flows, clearly exhibiting the coffee-stain effect, as depicted in panels (F-H). As evaporation continues and the reaction progresses, a secondary instability develops at the rim of the shrinking droplet, as shown in panel (H) (see Video 26).
2.2 Estimation of dimensionless numbers and timescale comparison
In our experiments, the chemical fingers forming on the pillars or on the walls of the PDMS/PMMA setups exhibit a wavelength of approximately $\lambda\approx 2~{}mm$. We hypothesize that the fluid fingers have a similar wavelength. Assuming the climbing fluid depth is in the range $h=20-100~{}\mu$[77, 78], we calculate the Marangoni stress $\tau$ using formula [69] $h=3\tau/2\rho g$, which represents the balance between the competing convective effects of Marangoni and gravitational forces. This calculation yields a surface stress range of $\tau=0.13-0.67$ $N/m^{2}$. Evaluating a solutal Marangoni number for $\tau=0.13$ $N/m^{2}$ and $h=20~{}\mu m$ yields:
$\mathscr{M}_{S}=\frac{\tau h^{2}}{\mu D}\sim\frac{0.13Nm^{-2}\times(20\times 10^{-6})^{2}m^{2}}{10^{-3}Nsm^{-2}\times 1.2\times 10^{-9}m^{2}s^{-1}}\sim 50,$
where $\mu$ is the viscosity of the CHD-BZ solution and $D$ is molecular diffusivity of chemical components. The thermal Marangoni number, defined as:
$\mathscr{M}_{T}=\frac{\tau h^{2}}{\mu\kappa},$
is two orders of magnitude smaller due to the ratio of molecular diffusivity to thermal diffusivity, given by
$D/\kappa=(1.2\times 10^{-9}m^{2}s^{-1})/(1.5\times 10^{-7}m^{2}s^{-1})\sim 10^{-2}.$
We now proceed by considering a dimensional analysis similar to Ref. [78] that examines the interaction between a deformable free surface and the steady-state concentration profile, which is governed by diffusion, evaporation, and convective transport along the pillars. Evaporation reduces the concentration of the more volatile chemical component with lower surface tension near the surface; however, under the assumption of a quasi-steady state, fresh fluid from the reservoir continuously replenishes the depleted region, establishing a stable vertical concentration gradient. We note that given the complexity of the CHD-BZ reaction, where multiple chemical components are continuously produced and consumed, the validity of this assumption requires further extensive investigation in future studies. If the free surface experiences an upward perturbation, the concentration profile must adjust before the interface returns to its original shape. If the concentration realigns more rapidly than the interface relaxes, the perturbed region develops a lower concentration than its surroundings, generating a surface tension gradient that draws fluid into the bump and amplifies the perturbation. Similarly, depressions in the free surface create local surface tension minima, pulling fluid away and reinforcing the deformation. This feedback loop sustains convection and enhances surface fluctuations, making the film susceptible to instability driven by Marangoni forces. The system’s stability depends on the relative timescales of diffusion and interface relaxation. If the diffusive timescale is significantly shorter than the relaxation time, the system is unstable, i.e., $T_{\mathrm{diff}}/T_{\mathrm{relax}}\ll 1.$ The relaxation time $T_{\mathrm{relax}}$ is estimated by balancing pressure gradients with viscous stresses,
$\nabla p\sim\left(\frac{\mu}{h^{2}}\right)V_{f}\sim\left(\frac{\mu}{h^{2}}\right)\frac{\lambda}{T_{relax}},$
where $V_{f}\sim 0.5$ $mm/s$ is the estimated speed of fluid movement up the pillars. The restoring pressure gradient is determined by either surface tension or gravity:
$p\sim\frac{\gamma h}{\lambda^{2}}\quad{\mathrm{or}}\quad p\sim\rho gh.$
For $h=20~{}\mu m$ and surface tension $\gamma=65~{}mN/m$ [79], the corresponding timescale comparison yields the dimensionless ratio for surface tension restoring pressure as:
$\frac{T_{\mathrm{diff}}}{T^{\gamma}_{\mathrm{relax}}}=\left(\frac{h}{\lambda}\right)^{4}\frac{\gamma h}{D\mu}\sim\left(\frac{20}{2000}\right)^{4}\frac{0.065\times 20\times 10^{-6}}{10^{-9}\times 10^{-3}}\sim 10^{-3},$
while for gravitational restoring pressure, it is
$\frac{T_{\mathrm{diff}}}{T^{g}_{\mathrm{relax}}}=\left(\frac{h}{\lambda}\right)^{2}\frac{h^{3}\rho g}{D\mu}\sim\left(\frac{20}{2000}\right)^{2}\frac{(20\times 10^{-6})^{3}\times 1000\times 10}{10^{-9}\times 10^{-3}}\sim 10^{-2}.$
Since both values are significantly less than one, we conclude that the system is susceptible to Marangoni-driven convection, where surface tension gradients enhance surface deformations and sustain instability. We note that accurately estimating the characteristic timescales and length scales in our study requires simultaneous experimental measurements of film thickness and concentration gradients driving Marangoni flows. Such simultaneous measurements are experimentally challenging, and efforts to address this are currently ongoing in our laboratory.
2.3 Numerical Simulations
Figure 12: Numerical simulations of the Tyson-Fife model (A) The space-time plot of concentrations $u$, $v$ and $w$ along the line marked in panel (B) illustrates that the obstacles become wave centers. (B) shows an image of the circular waves emanating from the obstacles at dimensionless time $t$$=$$7.5$. Time and space are in arbitrary units. (C) Simulations incorporating an asymmetric distribution of the inhibitor $w$ around the obstacles demonstrate wavefront instability at different time points. Only concentration of activator $u$ is shown (see Figure S6 for concentrations of both slow and fast inhibitors $v$ and $w$).
A full numerical simulation of the system would require solving the Navier-Stokes equations coupled with the chemical reaction dynamics. The Marangoni stress, imposed as a boundary condition on the deformed fluid’s free surface, connects fluid flow to chemical concentrations via surface tension gradients induced by concentration and thermal differences. While such simulations are beyond the scope of this manuscript and are currently being explored in our laboratory, we adopt a simplified approach that omits fluid flow and implicitly accounts for evaporation. Specifically, we assume that the reaction inhibitor (bromide ions) has a lower concentration near the pillars. Bromine gas, a product of the CHD-BZ reaction that mediates bromide ion production, is removed more rapidly in the vicinity of the obstacles, where the fluid ascends the hydrophilic pillars. As a result, the bromide ion concentration is effectively reduced on the pillars compared to the bulk solution. To explore the resulting wave dynamics, we employed the Tyson-Fife model [80] in two dimensions (see Materials and Methods for details). In our simulations, shown in Figure 12A-B, the inhibitor variable $w$ (bromide ions) was set to zero at the locations of the pillars. This localized depletion of the fast inhibitor $w$ near the obstacles leads to an increased concentration of the activator, which in turn facilitates the oxidation of the metal catalyst and the initiation of oxidation waves at the obstacles. These results align with our experimental observations.
We also analyzed the effect of an uneven angular distribution of $w$ around the pillars to emulate the disruption of wave fronts by Marangoni-driven fluid flows (refer to Materials and Methods for more information). The results of this analysis are presented in Figure 12C and Figure S6. Depending on the asymmetry introduced, a varying number of petals emerge following at least one initial cycle of circular waves. Remarkably, even with a set asymmetry of $w$ near the obstacles, the orientation of the petals can vary as the simulation progresses. Finally, to emulate the experimental setup of a tilted Petri dish (Figure S5A-B), we adapted the Tyson-Fife model to include a term that reflects a linear increase in fluid thickness (refer to Materials and Methods). In line with our experimental findings, the waves initially emerge in areas with thinner fluid and then move toward regions with thicker fluid (Figure S5C).
3 Conclusions
In this study, we introduce a periodic array of hydrophilic cylindrical obstacles into the CHD-BZ reaction, which rapidly become synchronized centers of wave activity. The thin layer of chemical solution ascending the obstacle surfaces is essential for enabling these structures to function as wave sources, thus precisely determining their spatial positions (Figure 1C). Although an isolated obstacle can also act as a wave center (Figure 7), an array arrangement effectively prevents the formation of random wave centers elsewhere in the solution. Under covered conditions, where evaporation is minimized, wavefronts maintain stable circular shapes through multiple cycles. In contrast, uncovered conditions—with significant evaporation—induce Marangoni-driven fingering instabilities, causing wavefront breakage and the emergence of striking flower-like patterns (Figure 1F). The number of petals in these patterns scales linearly with obstacle diameter (Figure 4), and there is a minimum diameter required for this instability to manifest (Figure S3). Additionally, control experiments where the weight of PDMS molds was monitored after each five-minute wave recording show negligible weight changes, conclusively ruling out any influence of PDMS swelling on our experimental results (Figure S8).
The behavior of a film climbing the exterior of a vertical or an inclined wall, influenced by the opposing forces of a thermally/chemically driven surface tension gradient and gravity, is extensively explored in the literature through numerical analysis using a thin-film model (see review articles [51, 52, 53] and references therein). In the case of a cylindrical obstacle, these models depend on three parameters: the cylinder radius, the upstream film thickness, and the downstream precursor film thickness. As the cylinder radius approaches infinity, the model converges to the scenario of a Marangoni-driven film ascending a vertical plate [54, 55, 56, 78]. Ref. [57] investigates the stability of the contact line by introducing an infinitesimal two-dimensional perturbation to the traveling wave solution, characterized by sinusoidal variations in the azimuthal direction. Through linear stability analysis, the study determines that the traveling wave (the so called Lax shock) is unstable to long-wave perturbations at the contact line. Given the periodic nature of the contact line in the azimuthal direction, only an integer number of fingers can form along it. The analysis identifies the most unstable mode and a neutrally stable cutoff mode, beyond which all modes are stable. These modes increase linearly with the cylinder radius, indicating that the number of fingers formed along the contact line correlates with the cylinder’s circumference. There exists a critical cylinder radius below which the contact line remains stable, with no fingers forming. Our experimental results provide direct validation of these theoretical predictions. In the future, we aim to expand the numerical analysis from Ref.[57] to include the rounded rectangular obstacle geometry (Figure 8), incorporating an asymmetric evaporation rate that results in an asymmetric surface tension gradient. Additionally, we will integrate the thin film equations from Ref. [57] with the reaction-diffusion equations to fully capture the dynamics of both the fluid flow and chemical components in the thin film surrounding the obstacles.
In our experiments, by tilting the Petri dish, it is possible to selectively control which array of obstacles initially becomes the focal point for wave generation (Figure S5A-B). As reported in Ref.[31] and consistent with our experiments, in a tilted Petri dish, waves originate in the thinner areas of the fluid, where convective flows are negligible, and progress toward the thicker regions, where convective flows can develop due to evaporative cooling. They assumed that these convective flows generate a heterogeneous medium where the refractory tail of a preceding wave annihilates part of the following wavefront, resulting in segmented waves that curl inward to form spirals. Over time, a chaotic wave pattern develops. In another study to understand the role of the fluid thickness in an uncovered Petri dish, Rossi et al.[40] have demonstrated that for fluid layers thinner than 1.0 mm, the BZ system behaves predominantly as a pure reaction-diffusion system, and convective effects are negligible. In layers with an intermediate thickness of 1.0–3.0 mm (the fluid thickness in our experiments was about 1.5 mm), Marangoni instabilities dominate, allowing rippled waves to propagate without breaking, and segmented patterns are not observed. However, for layers thicker than 3.0 mm, buoyancy-driven flows become significant, and wave segmentation processes occur. Our experiment using an uncovered, plasma-treated glass Petri dish with a fluid thickness of 2 mm also revealed rippled wavefronts and Marangoni-driven mosaic patterns (Figure S7A and Video 23) [42, 43, 44, 45]. We notice that at the boundary of the Petri dish, we observed a Marangoni-driven fingering instability, similar to the phenomena we see near the hydrophilic obstacles.
Contact angle measurements on both tapered and non-tapered pillars confirm that the contact line remains pinned (Figure 6C-E and 9C,H), indicating the presence of compensatory upward capillary flows. Additionally, experiments with tapered and inverse-tapered pillars demonstrate that as long as a wetting fluid layer surrounds the obstacle, it continues to function as a wave center. In tapered pillars, we observed that wavefront instability initially appears in straight pillars and, after a time delay, in pillars tapered at approximately 26. This delay is attributed to the higher evaporation rate in straight pillars due to their greater exposed surface area at the contact line. Interestingly, in a separate experiment investigating the effect of surface roughness, we 3D-printed conical pillars with a wavy surface (Figure S9). The results were largely consistent with our standard experiments without surface roughness, except for increased variability in petal size.
Finally, in our experiments with an evaporating CHD-BZ droplet, we observed the formation of spots at the rim. To elucidate the underlying mechanism, identify the more volatile components, and determine which chemical elements are segregated at the boundary, additional comprehensive experiments are necessary. Tracer beads added to the CHD-BZ solution in our evaporating droplet experiment reveal a pinned contact line and the coffee-stain effect, further suggesting the presence of compensatory capillary flows. Interestingly, after approximately one hour, we observe a complex secondary instability at the droplet boundary, which warrants further investigation. Ongoing studies in our laboratory are focused on exploring this phenomenon in greater detail.
In summary, our study on the spatio-temporal dynamics of chemical waves interacting with a periodic array of obstacles highlights several potential applications. By demonstrating how these obstacles can synchronize and guide wave propagation, we introduce a strategy for engineering specific wave patterns that could be especially beneficial in advanced microfluidic devices. We further propose that these guided patterns can be harnessed not only for exploring fundamental reaction–diffusion dynamics, but also for practical applications such as chemical computing, where controlled wave interactions serve as the basis for logic operations and information processing [81]. Moreover, our findings on manipulating Marangoni-driven flows around hydrophilic obstacles to shape wave patterns can be extended to colloidal droplet evaporation. By employing similar obstacle-based or surface-modification techniques, one could direct evaporative flows and steer colloidal self-assembly, opening new possibilities for creating uniform, tunable deposition patterns in printing, photonics, and related applications [82]. Finally, by adjusting parameters like obstacle geometry or evaporation conditions, one could similarly guide chemical wave propagation within gels, thereby influencing their mechanical oscillations. This capability paves the way for designing advanced biomimetic actuators or “smart” materials where desired wave-driven shape changes (or gel motions) can be precisely programmed or directed [83]. Overall, our results provide a foundation for further manipulation of chemical wave behaviors, offering exciting opportunities for technology and manufacturing [21, 84].
4 Method Section
[…]
Chemical preparation
The following chemicals and materials were used in this study. Sodium Bromate (NaBrO3), Sodium Bromide (NaBr), Conc. Sulfuric acid (Conc. H2SO4), 1,4-Cyclohexanedione, Ferrous sulfate, 1,10-phenanthroline was provided by sigma and used without any further purification. Ferroin was synthesized by mixing Ferrous sulfate hexahydrate and 1,10-phenanthroline in 1:3 molar ratio. For CHD-BZ experiments different concentration of the stock solutions were prepared in dH2O. Solution A: (494.58 mM solution of sodium bromate) Solution B: (891.9 mM solution of 1,4-Cyclohexanedione) Solution C: (971.9 mM solution of NaBr). A mixture was prepared in a beaker by combining 9 mL of solution A with 1.5 mL of solution B, which was then homogenized using a magnetic stirrer. Following this, 1.5 mL of the Ferroin indicator was added to the mix. Under a chemical fume hood, 0.8 mL of solution C was cautiously introduced to the reaction. The stirring continued for an estimated 1 hour and 45 minutes. Subsequent to this phase, an additional 1.5 mL of Ferroin was incorporated into the mixture. The stirring persisted at a controlled temperature of 23°C for a further 30 minutes, after which it was halted. Attention was then given to the mixture for any signs of color transformation, signaling the start of wave propagation, typically observed 3.5 to 4 hours post the initial mixing process. Following the onset of color change, the CHD-BZ reaction mixture was transferred to a plasma-treated PDMS mold featuring an array of macro-pillars, as specified in Figure 1A. These pillars, typically 1 mm in diameter and 3 mm high, are arrayed in a square grid with 5 mm spacing. In experiments with variable pillar dimensions and spacings, the volume of the CHD-BZ solution was adjusted to ensure a liquid height of about 1.5 mm, thus avoiding submergence of the pillars. Upon pouring, the solution was swirled gently until the mixture achieved a brown color. The Petri dish was then positioned undisturbed beneath a camera to record the wave propagation, under the illumination of white LED light. For each experiment, we record the waves for 5–10 minutes, observing until the waves from the Petri dish boundary interact with those centered around the obstacles. Afterward, we stop the recording, swirl the solution in the Petri dish to reset the system, and start recording again. Typically, the reagents are fully depleted after about 2 hours. In some experiments, a smaller, modified 50 mm Petri dish was used, and the volume of the CHD-BZ solution was correspondingly reduced to 1.5 mL to maintain the same liquid height. It’s important to note that in our standard setups, unless specified otherwise, the pillars do not submerge in the chemical solution. The hydrophilic nature of plasma-treated PDMS causes a thin layer of liquid to ascend the pillars (Figure 1B), altering the liquid surface near the obstacles as depicted in Figure S2C.
For some control experiments, the surface of the PDMS was coated with a 20 nm titanium layer and then with a pure gold layer. This was done using a Physical Vapor Deposition (PVD) technique with the Lesker Proline PVD 75 apparatus, rotating the substrate four times to achieve uniform coverage on the PDMS pillars’ surfaces.
[…]
Image Processing
The wave period is roughly 45 sec, and we capture images of the chemical waves at a rate of 1 frame per second across multiple wave cycles. To minimize spatial noise, we average the captured images over several wave cycles, typically 6-8, and then subtract each individual image from the time-averaged image. After this, we apply a Gaussian blur filter and use a customized MATLAB code to perform the Hilbert transform, which helps us calculate the phase map. This code is provided in the Supplementary Information.
[…]
Numerical simulations
We used the three-component Tyson-Fife model [80] to simulate our experimental system. The model focuses on the essential dynamics of the reaction by reducing the full chemical mechanism to a manageable number of reactions and variables. It typically includes the inhibitor $w$ which changes on the fastest time scale, activator $u$ which is the second fastest variable, and the slow varying inhibitor $v$. The variables $u$, $v$ and $w$ represent the concentrations of bromous acid (HBrO2), an oxidized form of the catalyst (such as ferroin Fe3+), and bromide ions, respectively. The set of equations for activator $u$ and two inhibitors $v$ and $w$ are:
$\frac{du}{dt}$ $=$ $D\nabla^{2}u+\frac{1}{\epsilon}(aw-uw+u-u^{2}),$ (1)
$\frac{dv}{dt}$ $=$ $u-v,$ (2)
$\frac{dw}{dt}$ $=$ $\frac{1}{\epsilon^{\prime}}(-aw-uw+bv),$ (3)
where $\epsilon^{\prime}\ll\epsilon\ll 1$, $a\ll 1$ and $b\simeq 1$. The parameters used in our simulations are based on those in Ref. [80]: $D=1$, ϵ=italic-$\epsilon=$0.05, ϵ=superscriptitalic-$\epsilon^{\prime}=$0.001, $b=$2.5, $a=$0.002 . We note that at $b=$2.5, the system is in the excitable state. The dimensions of the simulation box are chosen to be $N_{x}=N_{y}=1023$. The positions of the pillars on a square lattice are determined by $x$ and $y$ coordinates that meet the condition:$0.5[1+\cos(k(x-y)\cos(\pi/4))*\cos(k(x+y)\cos(\pi/4))]\leq 0.1$, where $k^{-1}$ representing the spacing between the pillars, is chosen to be $N_{x}/(6\pi)$. The concentration of inhibitor $w$ is set to zero inside the pillars.
In the simulations featuring an asymmetric distribution of $w$ around the pillars, we model the concentration of $w$ in a narrow ring around the pillars as: $w(x,y)=\alpha[1+\cos(k^{\prime}(x-y)\cos(\pi/4))*\cos(k^{\prime}(x+y)\cos(\pi/4))]$, where $k^{\prime}$ is arbitrarily set to $\pi/10$ and $\alpha=0.5$. After one cycle of circular waves, the wavefront breaks down and produces four petals, with the size of each petal varying over time, as depicted in Figure 12C.
In the tilted Petri dish experiment shown in Figure S5A-B, we adjusted the model to include a term proportional to the height $h(x)=x/N_{x}$ which varies linearly with $x$, under the assumption that the fluid height changes solely in the $x$-direction:
$\frac{du}{dt}$ $=$ $D\nabla^{2}u+\frac{1}{\epsilon}(aw-uw+u-u^{2}+p_{1}h(x)),$ (4)
$\frac{dv}{dt}$ $=$ $u-v+p_{2}h(x),$ (5)
$\frac{dw}{dt}$ $=$ $\frac{1}{\epsilon^{\prime}}(-aw-uw+bv+p_{3}h(x)),$ (6)
where $p_{1}=p_{2}=p_{3}=0.01$.
Supporting Information
Supporting Information is available from the Wiley Online Library or from the author.
Acknowledgements
We deeply appreciate the valuable insights and discussions provided by Professors G. Shubeita, A. Narayanan, A. Rebane, A. Pumir, O. Omelchenko, E. Frey, and D. Lohse. We also thank Dr. I. Gholami and M. Ibrahim for technical assistance and to the anonymous reviewers for their thoughtful comments and suggestions. We are especially grateful to the Core Technology Platform (CTP) at NYUAD for their exceptional support, with special thanks to Dr. R. Rezgui for his critical assistance with microscopy, Dr. Q. Zhang for his support in the clean room, and V. Dhanvi and J. Govindan for their outstanding contributions to fabricating the acrylic molds. Finally, we acknowledge the High Performance Computing (HPC) facility at NYUAD for providing essential resources for data storage and analysis, which were integral to this study.
Author contributions: S.G. and B.P. conducted the experiments, O.S. and A.G. designed and carried out the numerical simulations. A.G. conceived the idea, designed the experiments, analyzed the data, and wrote the first draft of the manuscript. All authors discussed the results and revised the manuscript.
Data and materials availability All data needed to evaluate the conclusions in the paper are present in the paper and/or the Supplementary Materials. Additional data related to this paper may be requested from the authors.
S5 Supplemental Figures
Figure S1: Frequency Analysis. (A) Color-coded period of the waves centered around the obstacles, measured in seconds. The Petri dish is covered and the waves maintain their circular shapes over multiple cycles, as shown in Figure 1K. (B) Averaged light intensity oscillations around three selected pillars highlighted in panel A, showing oscillations with periods of 45 seconds (magenta), 52.5 seconds (black), and 39.4 seconds (cyan). (C) Histogram of oscillations period. (D) Power spectrum of the oscillations shown in magenta in panel (B). (E-H) Similar to panels (A-D), but with the Petri dish uncovered, affecting the wave dynamics as shown in Figure 1L.
Figure S2: (A) A sample of a PDMS mold coated with a thin layer of gold is shown. The large pillars have a diameter of 3.8 mm. (B) Synchronous waves are centered around the gold-coated pillars. (C) A striped pattern is projected onto the fluid surface, highlighting the deformation caused by the rising fluid around the hydrophilic pillars. (D) A thin layer of hexane (with about 1 mm thickness) over the CHD-BZ solution prevents evaporation, ensuring the wavefronts preserve their circular shapes. Please note that the obstacles are not submerged. (E) Space-time plot along the cyan circle in panel D illustrating that over time the wavefronts maintain their circular shape.
Figure S3: No fingering instability in small-diameter pillars. (A-D) In experiments conducted with uncovered Petri dishes featuring pillars of 0.5 mm diameter, there is no occurrence of wavefront instability. This observation indicates that a certain minimum perimeter of the pillars is necessary for fingering instability to occur. The space-time plot shown in panel (D) is created by accumulating the light intensity around the cyan-colored circle depicted in panel (A).
Figure S4: Experiment in a glass Petri dish without obstacles. (A) A top view of a glass uncovered Petri dish with plasma treatment, illustrating fingering instability at the glass boundary. Notice that rippled waves travel and make visible the underlying mosaic patterns generated by the Marangoni instabilities. (B) A space-time plot along the cyan-colored semicircle shows the dynamics of the fingering, emphasizing that the positions of the fingers can change over time. (C-D) A setup similar to panel (A), but covered to minimize evaporation. In this configuration, wave centers appear at the periphery, but the wavefronts maintain their circular forms.
Figure S5: Phase map of an experiment in a tilted Petri dish (A) In a Petri dish with a slight tilt, synchronization waves initiate in the thinner part of the fluid and propagate towards the area with greater fluid thickness (video 15). (B) Space-time plot of the phase map along the green dashed line indicated in panel (A) illustrates the propagation of phase waves. Notice that in this experiment, the Petri dish is not covered, and flower patterns emerge at $t\sim 80$ sec. (C) Numerical simulations emulating the experimental setup shown in panel (A).
Figure S6: Numerical simulations with asymmetric $w$ around the pillar. (A) The space-time plot along the line marked in panel (B) illustrates that the obstacles become wave centers. (B) shows an image of the circular waves emanating from the obstacles. (C-D) Over time, the wavefront breaks, leading to the emergence of flower-like wave patterns.
Figure S7: Marangoni-driven surface flows. (A-B) In this experiment, tracer particles (20 $\mu$m in diameter) are used to visualize surface flows driven by Marangoni flows. These flows arise from chemical concentration gradients and slight temperature differences caused by evaporative cooling, which generate surface tension gradients (refer to video 27).
Figure S8: Negligible swelling of PDMS: To rule out PDMS swelling during our experiments, we measured the weight of the PDMS after each recording and found negligible changes. The initial PDMS weight at time zero is denoted as $W_{0}$. Data are presented for different PDMS-to-cross-linker ratios. All experiments were conducted using a PDMS-to-cross-linker ratio of 10:1. Even with a softer PDMS ratio of 10:0.5, the swelling effect remained negligible.
Figure S9: Experiments with Rough Obstacles: To examine the impact of surface roughness, we used a conical array of 3D-printed PLA (polylactic acid) obstacles with a wavy surface. Despite their rough texture, these obstacles continued to act as wave centers. In an open setup, the circular wavefronts fragmented, forming flower-like patterns, though with greater variability in petal size. The image in panel B is a processed version of the image in panel A, where the temporal mean has been subtracted, and a Gaussian blur has been applied.